OpenStudy (anonymous):

Assignment: Write the equation of the line and draw the graph of the line through each pair of points given. Use either the slope-intercept equation or by using x- and y-intercepts. (4, 6) and (-1, 2) (3, 5) and (3, 8) (5, 6) and (2, 6) (2, -1) and (0, 4) (you don't have to draw the graph I'm just confused as to how to write the equations

OpenStudy (jpes2193):

okay. start by finding m. then take you answer for m and plug it into y1-y=m(x-x1) and that should give your equation. i think this should help... let me know if this makes sense.

OpenStudy (anonymous):

@jpes2193 so would the first one be \[y = \frac{ 4 }{ 5 }x + \frac{ 14 }{ 5 }\]

OpenStudy (amorfide):